How do you write and solve equations in slope-intercept form for the line that satisfies the following condition: Passes through (6,11) and is parallel to the line that passes through (2,4) and (23,23) ?
Answer by mananth(16946)

x1 y1 x2 y2
2 4 23 23

slope m = (y2-y1)/(x2-x1)
( 23 - 4 )/( 23 - 2 )
( 19 / 21 )
m= 19/21
Parallel line will have same slope
slope = 19/21 and passing through (6,11)
Plug value of the slope and point ( 6 , 11 ) in
Y = m x + b
11.00 = 38/7 + b
b= 11 - 38/7
b= 5 4/ 7
So the equation will be
Y = 19/21 x + 39/7
m.ananth@hotmail.ca